OpenStudy (anonymous):

24x4y -------- 3x-4y-1 Which of the following is the simplest form of the expression above? Select one of the options below as your answer: A. 8 B. 8x^-16y^-1 C. 8x^-1y^-1 D. 8x^8y^2

OpenStudy (wolf1728):

24x4y -------- 3x-4y-1 Is that equation typed correctly? Shouldn't the numerator be something like 24x + 4y or 24x -4y?

OpenStudy (jack1):

none of the answers work tried reverse engineering the numerator from each answer times denominator... nothin

OpenStudy (jack1):

got it! equation is: ( 24 * x^4 * y) ------------- (3 * x^-4 * y^-1) \[\frac{ 24x^4 y }{3x^{-4}y^{-1}}\]

OpenStudy (jack1):

so split into like terms; equals: 24/3 * x^4 / x^-4 * y / y^-1 \[\frac{ 24 }{ 3 } \times \frac{ x^4 }{ x^{-4} } \times \frac{ y }{ y^{-1} } \] but \[\frac{ x^4 }{ x^{-4} } = \frac{ x^4 \times x^4 }{ 1 } \] and \[\frac{ y }{ y^{-1} } = \frac{ y \times y}{ 1 }\] so reduces to: \[\frac{ 24 }{ 3 } \times x^4 \times x^4 \times { {y} \times y} \] can you solve from here @8lue8lood ?

OpenStudy (cggurumanjunath):

d